the instructions state: Solve the Quatractic equation by factoring and using the Zero-Product Property.
the problem is: 3y^2-11y=10
here is what i've done so far:
i subtracted 10 from the right side so it will equal 0 then i added it to the binomial on the left.
i got 3y^2-11y+10=0
next, i would have to factor the trinomial. i got this far:
(3y^2-y)(-10y+10)
this is where i got stuck because i cant factor it correctly in order to use the Zero-Product Property.
Thank you for your help

Answer by richwmiller(17219)

You submitted this problem at least twice in one minute.
In the quadratic equation category yet you spelled Quatractic.
Do you need glasses? Seriously if you can't copy the word from right in front of you perhaps you need glasses.
The original equation is 3y^2-11y=10
"i subtracted 10 from the right side so it will equal 0 then i added it to the binomial on the left."
Didn't your teacher teach you that whatever you do on one side you have to do on the other? Yes the teacher did.
Think of the equation as a scale or balance or seesaw
You have to keep them the same.
3y^2-11y-10=0
It is funny that this equation can't be factored but 3y^2-11y+10=0 can be.
So double check your problem